# Towards Anytime Optical Flow Estimation with Event Cameras

### v1

@@ Abstract
Event cameras are capable of responding to log-brightness changes in microseconds. Its characteristic of producing responses only to the changing region is particularly suitable for optical flow estimation. In contrast to the super low-latency response speed of event cameras, existing datasets collected via event cameras, however, only provide limited frame rate optical flow ground truth, (e.g., at $10Hz$), greatly restricting the potential of event-driven optical flow. To address this challenge, we put forward a high-frame-rate, low-latency event representation Unified Voxel Grid, sequentially fed into the network bin by bin. We then propose EVA-Flow, an EVent-based Anytime Flow estimation network to produce high-frame-rate event optical flow with only low-frame-rate optical flow ground truth for supervision. The key component of our EVA-Flow is the stacked Spatiotemporal Motion Refinement (SMR) module, which predicts temporally-dense optical flow and enhances the accuracy via spatial-temporal motion refinement. The time-dense feature warping utilized in the SMR module provides implicit supervision for the intermediate optical flow. Additionally, we introduce the Rectified Flow Warp Loss (RFWL) for the unsupervised evaluation of intermediate optical flow in the absence of ground truth. This is, to the best of our knowledge, the first work focusing on anytime optical flow estimation via event cameras. A comprehensive variety of experiments on MVSEC, DESC, and our EVA-FlowSet demonstrates that EVA-Flow achieves competitive performance, super-low-latency ($5ms$), fastest inference ($9.2ms$), time-dense motion estimation ($200Hz$), and strong generalization. Our code will be available at https://github.com/Yaozhuwa/EVA-Flow.
Index Terms:
@@
I Introduction
Optical flow estimation is a fundamental task in computer vision, which aims to calculate the motion vector of each pixel in two consecutive frames of images. This dense prediction task has a wide range of applications in various fields, such as video compression [1], video frame interpolation [2], autonomous driving [3], robot navigation [4], etc. The accuracy of optical flow estimation directly affects the accuracy of subsequent autonomous perception tasks, making it a hot topic in the field of computer vision [5, 6, 7, 8, 9, 10].
Figure 1: (a-b) EVA-Flow vs. previous event-based optical flow networks. Our Event-based Anytime optical Flow (EVA-Flow) framework achieves time-dense optical flow estimation with low latency, making it especially suitable for high-speed event cameras and safety-critical scenarios. (c) Comparison between our dense motion trajectory and other methods. (d) EVA-Flow’s dense motion trajectory vs. E-RAFT [11] on the DSEC dataset [12].
@@ I Introduction
The event camera [13] is a new type of bio-inspired sensor that only responds to changes in the brightness of the environment. Compared with the traditional frame-based camera, which integrates the brightness at a certain time interval (exposure time) and outputs an image, the event camera has no exposure time and responds once the log value of the pixel intensity changes beyond a certain threshold, and the response is microsecond-level and asynchronous. As optical flow estimation aims to estimate the motion of a scene which aligns with the event camera’s ability to detect changes, it makes the event camera a suitable choice for optical flow estimation [14, 11]. Furthermore, event cameras offer distinct advantages for optical flow estimation, such as high temporal resolution and high dynamic range (${>}120dB$ compared to about $60dB$ of traditional frame-based cameras [13]). The high dynamic range capability of event cameras can solve the problem of under- or over-exposed images produced by traditional frame cameras in poor illumination conditions (such as at night) or high-dynamic scenes (such as cars entering and exiting tunnels) [15, 16], which can cause inaccurate optical flow estimation in subsequent processing. The high temporal resolution of event cameras eliminates the problem of motion blur even in high-speed motion scenes. Additionally, the high temporal resolution output of event cameras provides hardware support for super high-frame-rate optical flow estimation.
@@
However, most current event-based optical flow estimation methods fail to fully exploit the high temporal resolution and low-latency characteristics of events, resulting in low frame rates for the estimated optical flow. They typically transform the event stream over a time interval into a tensor represented as voxels, which are then fed into a frame-based optical flow estimation network similar to image-based approaches [14, 17, 11]. Moreover, the output frame rate is limited by the underlying dataset. The two most commonly-used benchmarks for event-based optical flow estimation, MVSEC [18] and DSEC [12] have low frame rates for their ground-truth optical flow, with DSEC having a frame rate of $10Hz$ and MVSEC having a frame rate of $20Hz$. While significant progress has been witnessed in the field, the limited temporal resolution greatly restricts the potential of event-based optical flow estimation in real-world applications.
@@
We present EVA-Flow, a deep architecture designed for Event-based Anytime optical Flow estimation. EVA-Flow delivers numerous benefits, such as low latency, high frame rate, and high accuracy. Fig. 1. (a-c) illustrates the differences between previous event-based optical flow estimation and our proposed approach toward anytime optical flow estimation. While other methods are limited by the dataset’s frame rate, this framework can achieve low latency and higher frame rate optical flow outputs. Fig. 1. (d) illustrates the discrepancy between the predictions of this model and E-RAFT using a sample from the test sequence of the DSEC dataset. The scene depicts a car making a turn inside a tunnel. Our EVA-Flow offers continuous trajectory tracking within a 100 ms time range of the input sample. The transition from yellow to red indicates varying point positions at different times in the trajectory of the car turning inside the tunnel. In contrast, E-RAFT [11] is restricted to generating a single optical flow.
@@
To achieve high-frame-rate event output, we first propose the Unified Voxel Grid (UVG), which generates a bin of UVG representation as soon as the events in a small time interval are ready. Our UVG representation achieves a higher frame rate up to a factor of $N$ and a lower latency of $1/N$ compared to the voxel grid representation [17]. $N$ is a hyperparameter that can be adjusted to meet the specific requirements of the task. These bins are then sequentially fed into our EVA-Flow, which consists of a multi-scale encoder and our stacked Spatiotemporal Motion Recurrent (SMR) module. SMR is proposed to predict temporally-dense optical flow and enhances the accuracy via spatial-temporal motion refinement. Notably, the architecture is specially designed for extreme low-latency optical flow estimation: as each event bin arrives, it is immediately processed through the SMR module, yielding the optical flow result for the current time instance. This architectural design eliminates the need to wait for all bins of the entire voxel grid to arrive in order to achieve high-frame-rate optical flow output.
Aside from the representation, for optical flow estimation, it is crucial to have a reliable method of measuring the consistency of the flow from one frame to the next. To achieve this, a loss function is usually designed to ensure that the optical flow converges. However, in the absence of ground truth for supervision, we do not directly supervise the intermediate optical flow, but instead, implicitly supervise it using the warping mechanism. Additionally, we introduce the Rectified Flow Warp Loss (RFWL) as a method to accurately assess the precision of event-based optical flow in an unsupervised manner. Moreover, we utilize RFWL to assess the accuracy of our time-dense optical flow estimation.
@@
By comparing our EVA-Flow with other event-based optical flow estimation methods on widely recognized public benchmarks, including the DSEC [12] and MVSEC [18] datasets, we have determined that our approach attains performance comparable to the state-of-the-art E-RAFT [11] method, with the added benefits of lower latency ($5ms$ vs. $100ms$), higher frame rates ($200Hz$ vs. $10Hz$), and faster single-frame optical flow estimation speed ($9.2ms$ vs. $93ms$). Additionally, the evaluation of time-dense optical flow on DSEC, MVSEC, and EVA-FlowSet underscores the reliability of our framework’s time-dense optical flow, even under the constraint of being solely supervised with low-frame-rate optical flow ground truth during training. Furthermore, the Zero-Shot results obtained from both the MVSEC dataset and our EVA-FlowSet clearly illustrate that our method exhibits superior generalization performance in comparison to E-RAFT.
In summary, we deliver the following contributions:
  • We present EVA-Flow, an Event-based Anytime optical Flow estimation framework that achieves high-frame-rate event-based optical flow with low-frame-rate optical flow ground truth as its sole source of supervision. EVA-Flow is characterized by its low latency, high frame rate, fast inference, and high accuracy.
  • We propose the Unified Voxel Gird representation, which is of high frame rate and low latency.
  • We propose RFWL to accurately evaluate the precision of event-based optical flow in an unsupervised manner.
  • Extensive experiments on public benchmarks show EVA-Flow achieves competitive accuracy, super-low latency, fastest inference, time-dense motion estimation, and strong generalization.
II Related Work
II-A Optical Flow Estimation
@@
Recently, the field of optical flow estimation has witnessed remarkable advancements owing to the progress of deep learning. FlowNet [19] firstly introduces an end-to-end network with a U-Net architecture to estimate optical flow directly from image frames and many subsequent works [20, 21, 22] follow this architecture. Later, [23, 24] utilize a warp mechanism to conduct a multi-scale refinement, achieving higher accuracy with less computational costs. RAFT [25] proposes a recurrent optical flow network that uses all-pair correlations and GRU iterations to produce high-precision flow results through iterative refinement, which achieves a significant improvement in accuracy. Most of the later optical flow studies are, in essence, improvements upon RAFT. CSFlow [26] introduces a decoupled strip correlation layer to enhance the capacity to encode global context. FlowFormer [27] explores the self-attention mechanism in recurrent flow networks to achieve further flow accuracy boosts, albeit with larger parameter usage. In summary, the use of warp mechanism, correlation volume, and RNN with iterative refinement has brought about significant advancements in the field of optical flow, allowing for improved accuracy and more efficient computation.
Unlike previous works, we explore anytime optical flow estimation with event cameras. We aim to overcome the limited time interval imposed by existing flow datasets and achieve ultra-high-frame-rate flow estimation with high accuracy.
II-B Event-based Optical Flow
@@
Event optical flow estimation is mainly divided into model-based methods and learning-based methods. There are two main approaches to model-based event optical flow estimation. One approach is to fit the local spatiotemporal plane of the event point cloud and use the plane fitting parameters to compute the optical flow value [28, 29, 30]. Another approach builds on the contrast maximization framework [31], which calculates the optical flow by optimizing the contrast of the motion-compensated event frame [32]. However, model-based event optical flow algorithms [33, 28, 29, 30] generally require denoising algorithms like [34] for event preprocessing to improve the accuracy.
@@
Thanks to the availability of large-scale event optical flow datasets [18, 12], learning-based event optical flow estimation algorithms [14, 17, 11] have achieved superior accuracy compared to model-based algorithms, which is also more robust to event noise. Next, we provide a brief review of works on learning-based event optical flow estimation.
EV-FlowNet [17] firstly utilizes bilinear interpolation on discrete events in both spatial and temporal dimensions. This transformation converts sparse and continuous events into a voxel-based tensor called Voxel Grid. Subsequently, the Voxel Grid is input to FlowNet [19] for estimating event optical flow. This paradigm has been adopted by numerous subsequent works [35, 11, 36, 37]. E-RAFT [11] achieves remarkable flow accuracy boosts by using event VoxelGrids with two-time intervals as the network’s input, recognized as the state of the art on the DSEC dataset [12]. Recently, TMA [38] and IDNet [39], have surpassed E-RAFT’s performance for their better utilization of the rich information in the temporal dimension of event data. The TMA network utilizes the temporal continuity of event input by calculating the correlation between the first bin and all other bins of the event Voxel Grid to acquire more precise motion information, achieving higher accuracy than E-RAFT with fewer iterations. IDNet sequentially feeds event Voxel Grids into RNN and estimates the flow for the entire time duration. Next, the network cascade and warp mechanisms conduct fine-grained processing on the flow, using the previously predicted flow in each cascade for direct deblurring of the event bins. While both TMA and IDNet make use of the continuous temporal information of event data, they can only achieve low-frame-rate optical flow outputs that match the frame rate of the ground truth, i.e., $10FPS$ on DSEC.
The event information provides brightness variation information of high temporal resolution, but only a few high-frame-rate event optical flow estimation methods are available. The limitation of the ground-truth frame rate of current datasets of event optical flow is one reason for this circumstance. Due to the lack of time-continuous ground truth, several studies [40, 41] have employed unsupervised contrast-maximization-based losses [32], resulting in lower accuracy for these methods. Ponghiran et al. [42] utilize recurrent neural networks to obtain temporally-dense optical flow results, which require sequential training and only evaluate the flow results at time locations with ground truth. The intermediate optical flow of this method has not undergone validation, and the final optical flow accuracy is relatively low. In contrast, our proposed approach, EVA-Flow, achieves a high level of accuracy, extremely low data latency, high frame rates, and excellent model generalization through comprehensive end-to-end training, requiring only low frame rate optical flow ground truth as supervision.
Figure 2: Overview of our proposed Event Anytime Flow Estimation (EVA-Flow) framework. The voxelized sequence input is processed sequentially by the encoder, and for each new event bin arrival, multi-scale feature maps are extracted. Based on these maps, the Spatiotemporal Motion Recurrent (SMR) module directly predicts optical flow for the time interval from $t_{0}$ to the current time step. $f^{i}$ denotes the features of the $i$-th level in a feature pyramid, while $f_{j}^{i}$ denotes the $j$-th event bin’s feature in $j^{i}$. $\mathbf{V}_{0,i}$ denotes flow prediction from time $t_{0}$ to $t_{i}$.
III Methodology

### v3

@@ Abstract
Event cameras respond to changes in log-brightness at the millisecond level, making them ideal for optical flow estimation. However, existing datasets from event cameras provide only low frame rate ground truth for optical flow, limiting the research potential of event-driven optical flow. To address this challenge, we introduce a low-latency event representation, Unified Voxel Grid, and propose EVA-Flow, an EVent-based Anytime Flow estimation network to produce high-frame-rate event optical flow with only low-frame-rate optical flow ground truth for supervision. Furthermore, we propose the Rectified Flow Warp Loss (RFWL) for the unsupervised assessment of intermediate optical flow. A comprehensive variety of experiments on MVSEC, DESC, and our EVA-FlowSet demonstrates that EVA-Flow achieves competitive performance, super-low-latency ($5\,ms$), time-dense motion estimation ($200Hz$), and strong generalization. Our code will be available at https://github.com/Yaozhuwa/EVA-Flow.
keywords:
@@
[…]
1 Introduction
@@
Optical flow estimation is a fundamental task in computer vision, which aims to calculate the motion vector of each pixel in two consecutive frames of images. The event camera Gallego et al. (2022) is a new type of bio-inspired sensor that only responds to changes in the brightness of the environment. Compared with the traditional frame-based camera, which integrates the brightness at a certain time interval (exposure time) and outputs an image, the event camera has no exposure time and responds once the log value of the pixel intensity changes beyond a certain threshold, and the response is microsecond-level and asynchronous. As optical flow estimation aims to estimate the motion of a scene which aligns with the event camera’s ability to detect changes, it makes the event camera a suitable choice for optical flow estimation Zhu et al. (2018); Gehrig et al. (2021). Furthermore, event cameras offer distinct advantages for optical flow estimation, such as high temporal resolution and high dynamic range (${>}120dB$ compared to about $60dB$ of traditional frame-based cameras Gallego et al. (2022)). The high dynamic range capability of event cameras can solve the problem of under- or over-exposed images produced by traditional frame-based cameras in poor illumination conditions (such as at night) or high-dynamic scenes (such as cars entering or exiting tunnels) Zhang et al. (2021, 2022); Zhou et al. (2024), which can cause inaccurate optical flow estimation in subsequent processing. The high temporal resolution of event cameras eliminates the problem of motion blur even in high-speed motion scenes. Additionally, the high temporal resolution output of event cameras provides hardware support for super high-frame-rate optical flow estimation.
Figure 1: (a-b) EVA-Flow vs. previous event-based optical flow networks. (c) Comparison between our dense motion trajectory and other methods. (d) EVA-Flow’s dense motion trajectory vs. E-RAFT Gehrig et al. (2021) on the DSEC dataset Gehrig et al. (2021).
@@
However, most current event-based optical flow estimation methods fail to fully exploit the high temporal resolution and low-latency characteristics of events, resulting in low frame rates for the estimated optical flow. They typically transform the event stream over a time interval into a tensor represented as voxels Zhu et al. (2018), which are then fed into a frame-based optical flow estimation network similar to image-based approaches Zhu et al. (2018, 2019); Gehrig et al. (2021). Moreover, the output frame rate is limited by the underlying dataset. The two most commonly-used benchmarks for event-based optical flow estimation, MVSEC Zhu et al. (2018) and DSEC Gehrig et al. (2021) have low frame rates for their ground-truth optical flow, with DSEC having a frame rate of $10Hz$ and MVSEC having a frame rate of $20Hz$. While significant progress has been witnessed in the field, the limited temporal resolution greatly restricts the potential of event-based optical flow estimation in real-world applications.
@@
We present EVA-Flow, a deep architecture designed for EVent-based Anytime optical Flow estimation. EVA-Flow delivers numerous benefits, such as low latency, high frame rate, and high accuracy. Fig. 1. (a-c) illustrates the differences between previous event-based optical flow estimation and our proposed approach toward anytime optical flow estimation. While other methods are limited by the dataset’s frame rate, this framework achieves ultra-low latency and offers optical flow estimation at a frequency 20 times higher. Fig. 1. (d) illustrates the discrepancy between the predictions of this model and E-RAFT Gehrig et al. (2021) using a sample from the test sequence of the DSEC dataset. The scene depicts a car making a turn inside a tunnel. Our EVA-Flow offers continuous trajectory tracking within a 100 ms time range of the input sample. The transition from yellow to red indicates varying point positions at different times in the trajectory of the car turning inside the tunnel. In contrast, E-RAFT is restricted to generating a single optical flow.
@@
To achieve high-frame-rate event output, we first propose the Unified Voxel Grid (UVG), which generates a bin of UVG representation as soon as the events in a small time interval are ready. Our UVG representation achieves a higher frame rate up to a factor of $N$ and a lower latency of $1/N$ compared to the voxel grid representation Zhu et al. (2019). $N$ is a hyperparameter that can be adjusted to meet the specific requirements of the task. These bins are then sequentially fed into our EVA-Flow, which consists of a multi-scale encoder and our stacked Spatiotemporal Motion Recurrent (SMR) module. SMR is proposed to predict temporally dense optical flow and enhances the accuracy via spatial-temporal motion refinement. Notably, the architecture is specially designed for extreme low-latency optical flow estimation: as each event bin arrives, it is immediately processed through the SMR module, yielding the optical flow result for the current time instance. This architectural design eliminates the need to wait for all bins of the entire voxel grid to arrive in order to achieve low-latency and high-frame-rate optical flow output.
Furthermore, we introduce the Rectified Flow Warp Loss (RFWL), a novel method for the precise, unsupervised evaluation of event-based optical flow precision, and apply it to assess the accuracy of our time-dense optical flow estimations.
@@
By comparing our EVA-Flow with other event-based optical flow estimation methods on widely recognized public benchmarks, including the DSEC Gehrig et al. (2021) and MVSEC Zhu et al. (2018) datasets, we have established that our approach achieves performance on par with leading methods, offering advantages such as reduced latency ($5\,ms$) and higher frame rates ($200Hz$). Additionally, the evaluation of time-dense optical flow on DSEC, MVSEC, and EVA-FlowSet underscores the reliability of our framework’s time-dense optical flow, even under the constraint of being solely supervised with low-frame-rate optical flow ground truth during training. Furthermore, the Zero-Shot results derived from both the MVSEC dataset and our EVA-FlowSet clearly illustrate the strong generalization capabilities of our method.
In summary, our key contributions are threefold:
  • EVA-Flow Framework: An EVent-based Anytime optical Flow estimation framework that achieves breakthroughs in latency and temporal resolution. The architecture features two novel components: (1) Unified Voxel Grid (UVG) representation enabling ultra-low latency data encoding, and (2) A time-dense feature warping mechanism where shared-weight SMR modules propagate flow predictions across temporal scales. This structural design fundamentally enables single-supervision learning - only final outputs require low-frame-rate supervision while implicitly regularizing intermediate timesteps through feature warping recursion.
  • Rectified Flow Warp Loss (RFWL): A new unsupervised metric specifically designed for evaluating event-based optical flow precision. This self-consistent measurement provides theoretical guarantees for temporal continuity validation of high-frequency optical flow estimation.
  • Systematic Validation: Comprehensive benchmarking demonstrates competitive accuracy, super-low latency, time-dense motion estimation and strong generalization capability. Quantitative analyses using RFWL further confirm the reliability of our continuous-time motion estimation.
2 Related Work
2.1 Optical Flow Estimation
@@
Recently, the field of optical flow estimation has witnessed remarkable advancements owing to the progress of deep learning. FlowNet Dosovitskiy et al. (2015) firstly introduces an end-to-end network with a U-Net architecture to estimate optical flow directly from image frames and many subsequent works Ilg et al. (2017); Hui et al. (2018, 2021) follow this architecture. Later, the works of Ranjan and Black (2017); Sun et al. (2018) utilize a warp mechanism to conduct a multi-scale refinement, achieving higher accuracy with less computational costs. RAFT Teed and Deng (2020) proposes a recurrent optical flow network that uses all-pair correlations and GRU iterations to produce high-precision flow results through iterative refinement, which achieves a significant improvement in accuracy. Most of the later optical flow studies are, in essence, improvements upon RAFT. CSFlow Shi et al. (2022) introduces a decoupled strip correlation layer to enhance the capacity to encode global context. FlowFormer Huang et al. (2022) explores the self-attention mechanism in recurrent flow networks to achieve further flow accuracy boosts, albeit with larger parameter usage. In summary, the use of warp mechanism, correlation volume, and RNN with iterative refinement has brought about significant advancements in the field of optical flow, allowing for improved accuracy and more efficient computation.
2.2 Event-based Optical Flow
@@
Event optical flow estimation can be categorized into two methodologies: model-based and learning-based approaches. The model-based approach to event optical flow estimation encompasses two distinct strategies. The first strategy involves fitting a local spatiotemporal plane to the event point cloud and subsequently leveraging the derived plane fitting parameters to ascertain the optical flow. This methodology has been extensively explored in the literature, as evidenced by the works of Benosman et al. Benosman et al. (2014), Mueggler et al. Mueggler et al. (2015), and Low et al. Low et al. (2020). The second strategy is predicated on the contrast maximization framework, as articulated by Gallego et al. Gallego et al. (2018). This approach entails the computation of optical flow by optimizing the contrast of a motion-compensated event frame, a process that has been further elucidated by Shiba et al. Shiba et al. (2022, 2024). However, model-based event optical flow algorithms Almatrafi et al. (2020); Benosman et al. (2014); Mueggler et al. (2015); Low et al. (2020) generally require denoising algorithms like Delbruck (2008) for event preprocessing to improve the accuracy.
@@
Thanks to the availability of large-scale event optical flow datasets Zhu et al. (2018); Gehrig et al. (2021), learning-based event optical flow estimation algorithms Zhu et al. (2018, 2019); Gehrig et al. (2021) have achieved superior accuracy compared to model-based algorithms, which is also more robust to event noise.
EV-FlowNet Zhu et al. (2019) firstly utilizes bilinear interpolation on discrete events in both spatial and temporal dimensions. This transformation converts sparse and continuous events into a voxel-based tensor called Voxel Grid. Subsequently, the Voxel Grid is input to FlowNet Dosovitskiy et al. (2015) for estimating event optical flow. This paradigm has been adopted by numerous subsequent works Ye et al. (2020); Gehrig et al. (2021); Sun et al. (2022); Hu et al. (2022); Lee et al. (2020). E-RAFT Gehrig et al. (2021), leveraging the sophisticated RAFT Teed and Deng (2020) optical flow estimation network, has markedly improved the precision of event-based optical flow estimation and has long been regarded as the state-of-the-art method on the DSEC-Flow dataset Gehrig et al. (2021). Recentlyy, E-Flowformer Li et al. (2023) leverages transformers Vaswani (2017) to enhance network performance, achieving higher optical flow estimation accuracy with an expanded synthetic dataset, BlinkFlow Li et al. (2023). TMA Liu et al. (2023) and IDNet Wu et al. (2024) have leveraged the abundant motion data recorded by event cameras during intermediate time frames to improve optical flow estimation, further enhancing the precision of event-based optical flow. Although these methods achieve high-precision optical flow estimation, they fail to capitalize fully on the low latency and swift response capabilities of event cameras. Constrained by the dataset’s frame rate, they are restricted to generating optical flow estimates at similarly low frame rates, specifically $10Hz$ on the DSEC-Flow dataset.
Event cameras offer brightness variation data with super-high temporal resolution; however, there is a paucity of methods for high-frame-rate event optical flow estimation. Due to the lack of time-continuous optical flow ground truth, several studies Chaney et al. (2021); Paredes-Vallés et al. (2023) have employed unsupervised contrast-maximization-based losses Shiba et al. (2022), resulting in lower accuracy for these methods. These self-supervised methods typically perform well on small optical flow datasets (MVSEC) but exhibit poor accuracy on datasets with large displacements (DSEC-Flow). Wan et al. Wan et al. (2022) develop an event-image fusion approach that combines events and images for temporal feature learning. While effective for continuous scene understanding, this method requires additional image inputs and recomputing event embeddings for each step during feature fusion. Ponghiran et al. Ponghiran et al. (2023) utilized recurrent neural networks to estimate time-dense optical flow from events, but their method required sequential training and was dependent on high-frame-rate ground truth optical flow for supervision. On the DSEC-Flow dataset, they employed interpolation to achieve high-frame-rate optical flow ground truth; however, due to the large displacements and the nonlinear nature of many motions within this dataset, their accuracy on DSEC-Flow was comparatively poor. Gehrig et al. Gehrig et al. (2024) introduced a continuous-time event optical flow algorithm that leverages dense temporal correlation to estimate pixel motion trajectory parameters (Bézier curves) across an entire time interval, facilitating continuous optical flow estimation. However, this method estimates the pixel motion trajectory for a specified time period only after the accumulation of all events ($100\,ms$ on DSEC-Flow). In contrast, our novel approach, EVA-Flow, attains high accuracy, minimal data latency, a very high prediction frequency, and superior model generalization through comprehensive end-to-end training, necessitating only low-frame-rate optical flow ground truth for supervision.
3 Method